Coves Cakes is a local bakery. Price and cost information follows: $ 17.00 Price per cake Variable cost per cake Ingredientsd. Sales price decreases by S0.50 per cake. Break-Even Point Bernak-even cakes 2. Assume that Cove sold 400 cakes last month.
Answer #1

1 (a)

Sales price = 17 + 1

= $18 per cake

Total Variable cost = 2.50 + 1.40 + 0.20

= $4.10 per cake

Contribution margin = Sales price - Variable cost

= 18 - 4.10

= $13.90 per cake

Break even point = Fixed cost/Contribution margin

= 3,850/13.90

= 277 cakes

1 (b)

Fixed cost = 3,850 + 500

= $4,350

Variable cost = 2.50 + 1.40 + 0.20

= $4.10 per cake

Contribution margin = Sales price - Variable cost

= 17 - 4.10

= $12.90 per cake

Break even point = Fixed cost/Contribution margin

= 4,350/12.90

= 337 cakes

1 (c)

Variable cost = 4.10 - 0.35

= $3.75 per cake

Contribution margin = Sales price - Variable cost

= 17 - 3.75

= $13.25 per cake

Break even point = Fixed cost/Contribution margin

= 3,850/13.25

= 291 cakes

1 (d)

Sales price = 17 - 0.50

= $16.50 per cake

Variable cost = 2.50 + 1.40 + 0.20

= $4.10 per cake

Contribution margin = Sales price - Variable cost

= 16.50 - 4.10

= $12.40 per cake

Break even point = Fixed cost/Contribution margin

= 3,850/12.40

= 310 cakes

2.

Income statement

Sales (400 x 17)

6,800

- Variable cost (400 x 4.10)

- 1,640

Contribution

5,160

- Fixed cost

- 3,850

EBIT

1,310

Degree of operating leverage = Contribution/EBIT

= 5,160/1,310

= 3.94

3.

Degree of operating leverage = % change in EBIT/% change in sales

3.94 = % increase in EBIT/10%

Hence, % increase in EBIT = 39.4%

Hope this helped ! Let me know in case of any queries.